OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

what are the answers

OpenStudy (anonymous):

ok 3x - y = -3 and passes through the point (4, -2). 1st you subtract the 3x from the side with the y and put it on the side with the -3

OpenStudy (anonymous):

y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

That's the answers

OpenStudy (anonymous):

It then is -y=-3x-3 so you divide the negative 1 from the y and get -1y/-1=-3x/-1 and -3/-1 to get

OpenStudy (anonymous):



OpenStudy (anonymous):

Is it negative?

OpenStudy (anonymous):

the answer is 3x-14=-2

OpenStudy (anonymous):

Also um, that's not an answer

OpenStudy (anonymous):

how not

OpenStudy (anonymous):

oh it isn't

OpenStudy (anonymous):

Unless it's y = 3x - 14

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i think it is as well

OpenStudy (anonymous):

I think thats the answer 3x-14

OpenStudy (anonymous):

okay thx